### v1

# Predictive Modeling of Flexible EHD Pumps using Kolmogorov-Arnold Networks

Abstract
We present a novel approach to predicting the pressure and flow rate of flexible electrohydrodynamic pumps using the Kolmogorov-Arnold Network. Inspired by the Kolmogorov-Arnold representation theorem, KAN replaces fixed activation functions with learnable spline-based activation functions, enabling it to approximate complex nonlinear functions more effectively than traditional models like Multi-Layer Perceptron and Random Forest. We evaluated KAN on a dataset of flexible EHD pump parameters and compared its performance against RF, and MLP models. KAN achieved superior predictive accuracy, with Mean Squared Errors of 12.186 and 0.012 for pressure and flow rate predictions, respectively. The symbolic formulas extracted from KAN provided insights into the nonlinear relationships between input parameters and pump performance. These findings demonstrate that KAN offers exceptional accuracy and interpretability, making it a promising alternative for predictive modeling in electrohydrodynamic pumping.
Index Terms:
Kolmogorov-Arnold Networks, Electrohydrodynamic pumps, Neural network.
I Introduction
The electrohydrodynamic (EHD) pumps are devices that harness electrostatic forces to induce the movement of a dielectric fluid, offering several advantages over traditional mechanical pumps, such as the absence of moving parts, high efficiency, and low noise [1]. EHD pumps have been studied extensively due to their unique ability to generate fluid flow without moving mechanical components. Traditional EHD pumps consist of a series of electrodes that induce fluid flow by ion drag, dielectrophoresis, or electrothermal effects. Flexible EHD pumps, a newer development, offer enhanced integration capabilities with soft actuators and biomedical devices due to their adaptable and lightweight nature. These characteristics make EHD pumps ideal for applications in microfluidics, cooling systems, soft robotics, and biomedical devices. A conventional EHD pump consists of electrodes placed within a fluid channel, creating an electric field that generates ions or dipoles. The movement of these charged particles results in fluid flow and pressure generation.
Recently, flexible EHD pumps have gained traction due to their adaptability and potential for integration into soft actuators and fluidic systems. However, predicting the performance of flexible EHD pumps remains a challenge due to the intricate interactions between electrical, mechanical, and fluidic fields. Traditional physical simulations often neglect fluid-solid interactions and energy losses associated with flexible materials, making accurate performance prediction difficult. Consequently, there is a growing interest in using machine learning models to predict critical parameters such as pressure, flow rate, and power.
In previous study [2], three machine learning models were employed—random forest (RF), ridge regression (RR), and multi-layer perceptron (MLP) to predict the performance of flexible EHD pumps. This study demonstrated that MLP, a type of feedforward artificial neural network, outperformed other models in predicting the pressure and flow rate of flexible EHD pumps. However, MLP typically rely on fixed activation functions at the nodes (or ”neurons”), which can constrain the model’s ability to learn complex and nonlinear relationships efficiently. Moreover, MLP can require a substantial number of neurons and layers to achieve high accuracy, leading to increased computational complexity.
Recently, Kolmogorov-Arnold Networks (KAN) was proposed to promise alternatives for MLP [3]. Inspired by the Kolmogorov-Arnold representation theorem, KAN replace fixed activation functions with learnable activation functions on the edges (”weights”). This architectural difference allows KAN to approximate nonlinear functions more effectively while maintaining a smaller network size compared to MLP. Despite their expressive power, MLP often require a substantial number of neurons and layers to achieve high accuracy, leading to overfitting risks and computational inefficiency. KAN address these limitations by incorporating learnable spline functions at each weight, providing greater flexibility and accuracy in learning nonlinear relationships. This unique architecture offers faster neural scaling laws and requires fewer parameters compared to MLP, as demonstrated in various mathematical and physics-based tasks.
This study developed and evaluated a KAN-based predictive model for the output performance (pressure and flow rate) of flexible EHD pumps and compared it against the MLP and Random Forest models. The contributions of this paper are introducing a novel application of KAN for predicting the performance of flexible EHD pumps and providing insights into the relations of input parameters in determining the output performance of flexible EHD pumps.
II Methodology
The Kolmogorov-Arnold representation theorem. The theorem states that any multivariate continuous function can be represented as a finite composition of continuous univariate functions and the binary operation of addition. More formally, for a smooth function $\textit{f}:[0,1]^{n}\to\mathbb{R}:$
$f(x)=f(x_{1},\ldots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right)$ (1)
Where $\phi_{q,p}:[0,1]\to\mathbb{R}$ and $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ are continuous functions. Essentially, the theorem implies that any high-dimensional function can be broken down into a sum of simpler univariate functions.
Although the Kolmogorov-Arnold theorem is theoretically powerful, earlier attempts to apply it in machine learning encountered challenges due to the fractal nature of some functions and the complexity of representation. However, recent advances in neural network architectures have revitalized interest in this approach. Liu et al. proposed the KAN, which leverages spline-based activation functions and a grid extension technique to achieve remarkable accuracy in function approximation [3]. The study compared KAN to MLP in various data fitting tasks and found that KAN could achieve comparable or better accuracy with fewer parameters. Specifically, a 2-layer KAN with a width of 10 was found to be 100 times more accurate than a 4-layer MLP with a width of 100 in solving partial differential equations (PDEs). Furthermore, KAN demonstrated the interpretability advantages. By visualizing the learned spline functions, researchers can gain insights into the underlying mathematical or physical laws captured by the model. This characteristic makes KAN particularly attractive for scientific discovery and collaboration.
Inspired by the mathematical representation of the Kolmogorov-Arnold theorem, function $f(x)$ is expressed as a composition of inner and outer function matrices applied to input vector ${\bf x}$, represented as:
$f(x)={\bf\Phi}_{\rm out}\circ{\bf\Phi}_{\rm in}\circ{\bf x}$
Here, ${\bf\Phi}_{\rm in}$ is a matrix of univariate functions, denoted as:
${\bf\Phi}_{\rm in}=\begin{pmatrix}\phi_{1,1}(\cdot)&\cdots&\phi_{1,n}(\cdot)\\ \vdots&&\vdots\\ \phi_{2n+1,1}(\cdot)&\cdots&\phi_{2n+1,n}(\cdot)\end{pmatrix}$
and ${\bf\Phi}_{\rm out}$ is a row vector of univariate functions:
${\bf\Phi}_{\rm out}=\begin{pmatrix}\Phi_{1}(\cdot)&\cdots&\Phi_{2n+1}(\cdot)\end{pmatrix}$
These matrices illustrate a Kolmogorov-Arnold layer, which forms the basis of the KAN by stacking such layers. A KAN is thus constructed as:
${\rm KAN}({\bf x})={\bf\Phi}_{L-1}\circ\cdots\circ{\bf\Phi}_{1}\circ{\bf\Phi}_{0}\circ{\bf x}$
This structure contrasts with the MLP, which alternates linear layers and nonlinear activation functions as follows:
${\rm MLP}({\bf x})={\bf W}_{L-1}\circ\sigma\circ\cdots\circ{\bf W}_{1}\circ\sigma\circ{\bf W}_{0}\circ{\bf x}$
KAN builds on this theoretical foundation by replacing fixed activation functions with learnable activation functions on the edges (weights). Unlike traditional MLP, KAN have no linear weight matrices. Instead, each weight is represented as a learnable spline function. This architectural innovation allows KAN to better capture complex, nonlinear relationships by optimizing univariate functions directly.
The visualization of KAN is straightforward, with each layer resembling a fully-connected layer where each edge holds a one-dimensional function, simplifying the understanding of complex networks through a direct representation of function layers. This model allows for a deeper and more interpretable connection between input features and outputs, showcasing the potential for enhanced model transparency in machine learning applications.
In this study, we utilized three machine learning models, MLP, RF and KAN, to predict the pressure and flow rate of flexible EHD pumps. Each model offers unique strengths and weaknesses, providing a comprehensive comparison.
MLP is a type of feedforward artificial neural network consisting of an input layer, multiple hidden layers, and an output layer. Each layer contains neurons that use fixed nonlinear activation functions $tanh$. In this study, we used an MLP model with three hidden layers and trained using the Adam optimizer with a learning rate of 0.001. RF is an ensemble learning method that constructs multiple decision trees during training and aggregates their results to improve predictive accuracy. In this study, we employed an RF model with 100 decision trees. The model’s maximum depth was set to 10 to prevent overfitting. The Gini impurity criterion was used to measure the quality of the splits. KAN replaces fixed activation functions in MLP with learnable univariate spline functions at the edges. We used two different KAN models:
  • For pressure prediction: KAN(width=[5, 2, 1], grid=2, k=3, seed=0)
  • For flow rate prediction: KAN(width=[5, 6, 1], grid=4, k=4, seed=0)
The dataset was divided into training and testing subsets using a 90-10 split, resulting in 88 training samples and 10 testing samples. For predicting pressure, we created a KAN model with two hidden layers (width=[5, 2, 1]) and cubic splines (k=3). The training was carried out using the LBFGS optimizer, which is suitable for small datasets. After training, the model was pruned to simplify its structure while maintaining its predictive accuracy. The pruned model was re-trained for further refinement. To increase interpretability, we used symbolic formula extraction to approximate the learned spline functions with mathematical expressions. For predicting flow rate, a KAN model with two hidden layers (width=[5, 6, 1]) and cubic splines (k=4) was used. The same pruning, refinement, and symbolic formula extraction steps were then repeated to simplify and interpret the flow rate prediction model.
III Experimental Setup
In this study, we used a series of flexible EHD pumps fabricated using microfluidic techniques to investigate their output performance in terms of pressure and flow rate. The flexible EHD pumps consist of polypropylene (PP) sandwiched electrodes, forming a fluidic channel that drives fluid through an applied electric field.
Each pump was connected to a fluid reservoir using flexible tubing. The entire system was evacuated to remove any gas bubbles and ensure consistent measurements. The fluid used in all experiments was Novec 7300, chosen for its thermal and chemical stability. The flexible EHD pumps were positioned at the same level as the fluid reservoir to obtain accurate low-pressure measurements, calculated using the formula $p=\rho gh$, where:
$\rho$ is the fluid density, $g$ is the gravitational acceleration, and $h$ is the height difference between the liquid surface in the reservoir and the EHD pump outlet. For precise pressure measurements, a pressure sensor (KEYENCE GP-M001) was employed. The generated flow rate was determined by leveling the liquid in the reservoir and the outlet tube to the same height and using a digital balance to measure the average flow rate. The current was measured using an ammeter, while voltage was applied through a DC high-voltage power supply (HEOPT-20B10, Matsusada Precision).
The critical geometric parameters of the flexible EHD pumps were varied systematically to obtain comprehensive data. These parameters include:
  • Channel Height: Three sizes were tested: 1 mm, 0.5 mm, and 0.15 mm.
  • Overlap of Electrode Pairs: Three overlaps were studied: 8 mm, 4 mm, and 0 mm.
  • Voltage: The voltage was set in the range of 0 to 11 kV.
  • Gap between Electrodes: Four gaps were measured: 0.3 mm, 0.6 mm, 0.9 mm, and 1.2 mm.
  • Apex Angle of Electrodes: Four angles were investigated: $\pi$, $\pi/2$, $\pi/3$, and $\pi/6$.
Each measurement was repeated three times to ensure accuracy, and the average values were used for further analysis.
The experiments generated a comprehensive dataset of 98 samples, each containing five input features and two output features. The input features are channel height, electrode overlap, voltage, electrode gap and apex angle, while the output features comprise the maximum pressure (Pa) and maximum flow rate (ml/min) of the flexible EHD pumps. These features are represented as: $\mathbf{X}=[X_{1},X_{2},X_{3},X_{4},X_{5}]$, where: $X_{1}$ is the channel height (mm), $X_{2}$ is the electrode overlap (mm), $X_{3}$ is the applied voltage (V), $X_{4}$ is the gap between electrodes (mm) and $X_{5}$ is the apex angle (°). The output vector $Y$ is defined as: $\mathbf{Y}=[Y_{1},Y_{2}]$, where: $Y_{1}$ is the maximum pressure (Pa), and $Y_{2}$ is the maximum flow rate (ml/min). The dataset was randomly split into training and testing subsets using a 90-10 split, resulting in 88 training samples and 10 testing samples. The training set was used to develop and optimize the machine learning models, while the testing set served to evaluate the models’ performance on unseen data. Mean Squared Error (MSE), which quantifies the average squared difference between predicted and actual values, was used to compare the performance of different machine learning models.
Figure 1: Parametrize splines in (a) pressure prediction and (b) flow rate prediction.
IV Result and Discussion
After training and evaluating all models for both pressure and flow rate prediction, the results were compared comprehensively. KAN showed significant improvements in both predictive accuracy and interpretability due to its learnable activation functions and spline-based architecture.
IV-A Pressure prediction
Predicting the pressure of flexible EHD pumps involves understanding complex nonlinear interactions among various input parameters such as voltage, channel height, apex angle, electrode overlap, and electrode gap. In this study, we used the KAN model to predict pressure. The KAN model utilized learnable spline-based activation functions to approximate nonlinear functions effectively. For pressure prediction, we configured KAN with two hidden layers (width=[5, 2, 1]) and cubic splines (k=3). The training process involved several stages, including:
1. Initial training with sparsity regularization.
2. Pruning to obtain a smaller model shape.
3. Refinement through further training and symbolic formula extraction.
Figure 1 (a) demonstrates the basis functions ($B_{i}(x)$) used by the KAN model for pressure prediction. Each function is parameterized as a B-spline, which adaptively learns the relationships between input features and output pressure.
Figure 2 (a) visualizes the different stages of training and pruning, demonstrating how the model was refined progressively. The training began by plotting the initial KAN structure, followed by training with sparsity regularization to simplify the model. Next, the KAN model was pruned to obtain a smaller shape, and training continued to refine the model further. Finally, symbolic formula extraction was used to approximate the learned spline functions with mathematical expressions.
KAN significantly outperformed other models in pressure prediction. The MSE was markedly lower, showcasing KAN’s superior predictive power and efficiency. The symbolic formula extracted from KAN provided further interpretability, allowing us to understand the underlying relationships between input features and pressure. KAN demonstrated exceptional predictive accuracy and interpretability, making it a promising alternative to traditional machine learning models in predicting the performance of flexible EHD pumps.
IV-B Flow rate prediction
For flow rate prediction, we configured the KAN model with two hidden layers (width=[5, 6, 1]) and cubic splines (k=4). The training process involved several stages was same with the pressure prediction.
Figure 1 (b) demonstrates the basis functions ($B_{i}(x)$) used by the KAN model for flow rate prediction. Each function is parameterized as a B-spline, adaptively learning the relationships between input features and output flow rate. These spline functions enable KAN to approximate complex nonlinear functions effectively while maintaining interpretability.
Figure 2 (b) visualizes the different stages of training and pruning, demonstrating how the model was refined progressively. The training began by plotting the initial KAN structure, followed by training with sparsity regularization to simplify the model. Next, the KAN model was pruned to obtain a smaller shape, and training continued to refine the model further. Finally, symbolic formula extraction was used to approximate the learned spline functions with mathematical expressions.
KAN significantly outperformed other models in flow rate prediction due to its ability to capture complex relationships between input features. The lower MSE, indicating an excellent fit between the predicted and actual flow rate values. The symbolic formula extracted from KAN provided further interpretability, enabling us to understand the underlying relationships between input features and flow rate.
Figure 2: Training process in (a) pressure prediction and (b) flow rate prediction.
IV-C Evaluation
KAN achieved superior predictive accuracy for pressure and flow rate prediction are depicted in Table I. The symbolic formula extracted from KAN provided further interpretability, allowing us to understand the underlying relationships between input features and pressure. The formula is as follows:
$\begin{split}Y_{1}&=12.46\cdot\exp(-0.01\cdot(1-0.8\cdot x_{2})^{3}\\ &+4.3\cdot(1-0.75\cdot x_{4})^{4}\\ &-0.06\cdot\exp(3.03\cdot x_{1})\\ &+3.18\cdot\tanh(0.18\cdot x_{3}-0.55)\\ &+2.43\cdot\exp(-2.57\cdot x_{5}))-1.87\end{split}$
Similarly, predicting the flow rate of flexible EHD pumps requires capturing complex nonlinear relationships between input features. The symbolic formula extracted from KAN provided further interpretability for flow rate prediction:
$\begin{split}Y_{2}&=1.7-1.59\cdot\tanh(22.4\cdot(0.9-x_{4})^{4}\\ &-3.33\cdot\sin(6.2\cdot x_{3}-2.35)\\ &+0.08-2.11\cdot\exp(-1.72\cdot x_{5})\\ &+2.13\cdot\exp(-0.24\cdot x_{2})\\ &-0.89\cdot\exp(-1.4\cdot x_{1}))\end{split}$
TABLE I: MSE for pressure and flow rate predictions by different models.
Model Pressure MSE Flow Rate MSE
KAN 12.186 0.012
Random Forest 1750.017 0.040
MLP 78.329 0.002
IV-D Discussion
The comparative analysis of the predictive models demonstrates that the KAN is superior in both pressure and flow rate predictions for flexible EHD pumps. KAN’s advantage lies in its unique architecture, which replaces fixed activation functions with learnable spline-based activation functions on the edges (”weights”). This architecture, inspired by the Kolmogorov-Arnold theorem, allows KAN to approximate nonlinear functions more effectively than traditional models. The KAN model exhibited remarkable predictive accuracy, with MSE values of 12.186 and 0.012 for pressure and flow rate predictions, respectively. These results significantly outperform Random Forest, and MLP models (in pressure prediction), highlighting KAN’s superior ability to capture complex nonlinear relationships. Furthermore, KAN’s symbolic formula extraction feature provides interpretability that traditional neural networks often lack. The symbolic formulas reveal the underlying mathematical relationships between input features and target variables, offering insights into how each feature affects the pressure and flow rate of the EHD pump.
The symbolic formulas extracted from KAN indicate the relative importance of input parameters. For instance, the pressure formula shows that the voltage ($x_{1}$), channel height ($x_{2}$), and electrode gap ($x_{5}$) have significant nonlinear influences on pressure. The flow rate formula demonstrates that apex angle ($x_{3}$), electrode overlap ($x_{4}$), and gap ($x_{5}$) are crucial determinants of flow rate. Such insights can guide the design and optimization of flexible EHD pumps.
Moreover, this study focused on flexible EHD pumps with a specific set of input features. Future work could investigate the generalizability of KAN to other types of pumps or predictive tasks involving different input parameters. The KAN offers a novel and effective approach to predicting the pressure and flow rate of flexible EHD pumps. Its unique architecture, predictive accuracy, and interpretability make it a valuable tool for researchers and engineers in the field of electrohydrodynamic pumping.
V Conclusion
In this study, we explored the predictive performance of the KAN in forecasting the pressure and flow rate of flexible EHD pumps. The KAN replaces fixed activation functions with learnable spline-based activation functions, enabling it to approximate complex nonlinear functions more effectively than traditional models. This network demonstrated superior predictive accuracy in pressure predictions compared to other models. For pressure prediction, KAN achieved a MSE of 12.186, while in flow rate prediction, KAN obtained an MSE of 0.012. These results showcased KAN’s exceptional ability to capture the intricate nonlinear relationships between input parameters and output performance metrics. Moreover, the interpretability of KAN through symbolic formula extraction provides valuable insights into the relationships between input features and output variables. The symbolic formulas reveal the significant nonlinear influence of parameters of voltage, apex angle, and electrode gap on the pressure and flow rate of EHD pumps. Such interpretability makes KAN a valuable tool not only for accurate predictions but also for guiding design and optimization in electrohydrodynamic pumping.